# The PLATO Mission

### v1

@@ 4.3.1 Expected solar-like oscillations within the PLATO P1-P2 sample
The first question is to what extent solar-like oscillations can be detected for PLATO targets. Figure 11 shows the stars in the P1–P2 sample for two long-pointing fields in an HR diagram. They are taken from the PIC version PICv1.1.0 [229], which provides effective temperatures and stellar radii, hence luminosities. In Fig. 11, we distinguish between stars for which a theoretical calculation has led to a probability of detection of solar-type oscillations and those for which the threshold for a positive seismic detection has not been reached. The probability of detection were not calculated for hot stars in the instability band defined according to the criterion adopted by [69]. This concerns few stars because the restriction on the temperature of the hot side adopted to construct the PIC is more severe. We also did not consider the early red giants –also contained in the PIC– which will be included in a specific scientific calibration sample catalogue as part of the overall PIC [5]. Their performance will be addressed elsewhere. Once early red giants and hot stars have been removed, the total subsample (dwarfs and subgiants) contains 14083 stars. We considered a positive seismic detection when the probability of the signal being due to noise is below 0.1% and the probability of detecting solar-like oscillations is equal to 99% or above. The computation of the probability was adapted from the one described in Chaplin et al. (2011a). The seismic detection probability depends on the observing duration. We therefore show two cases in Fig.11 : observing runs of 30 days and 2 years long. The probability also involves the signal-to-noise ratio in a power spectrum. We then used a combination of the formulations by [69] and by [268] for the oscillation amplitudes calibrated with Kepler data (a compilation of oscillating stars from the catalogues of [278] and [206] and short-cadence stars found with no detection from the [205] catalogue. For the noise, we used the PLATO (random and systematic residuals) noise level included in the PIC, to which we added the stellar granulation background noise. For comparison, we included in the plot the positions of some Kepler stars from the Legacy sample [198, 283] as characterised by [90]. The calculation of the probability to detect oscillations also depends on the width () of the assumed Gaussian-shape envelope due to oscillations in a power spectrum. From observations, ranges from to . The first option ( together with a probability threshold set at 0.90) provides a number of positive detections in agreement with the number of Kepler stars with detected oscillations, but a too large number of false-positive detections among the Kepler stars with no detected oscillations. For the other option ( together with a conservative probability threshold of 0.99), it is the reverse, i.e., the number of predicted non-detections is in agreement with the number of Kepler stars without oscillations, but too few detections are found compared to the Kepler stars with detected oscillations. Here we remain conservative and use the second option to derive the number of expected detections for PLATO. We count 11408 PLATO P1–P2 stars – including 6137 subgiants – for which detections of solar-like oscillations are expected after 2 years of observation. These stars span the whole range of mass and age of the P1–P2 sample. With 2 years of observation, the number of MS-stars with expected detection is about 5271. The impact of the observation duration is significant: from 30 days to 2 years and limiting the count to stars with masses , the number of detections increases from 324 ( 7.5% of stars with masses ) to 2393 ( 55% of stars with masses ). Had we assumed the other option, the number of expected detections for stars with mass would increase up to 3291 (76% of stars with masses ) for a 2 years run. Not surprisingly, stars for which one might not detect solar-like oscillations for too short an observing time are MS- stars of low mass because their oscillation amplitudes are too small. In contrast, detection of solar-like oscillations - when applied to the Kepler sample- is not obtained for some stars for which they are theoretically expected [69, 68, 205]. Several reasons may occur, one being the strong magnetic activity. As an example, we found 8 % false-positive detections for the [205] Kepler sample. Based on this, the number of PLATO detections for main-sequence stars with mass decreases to 51%.
@@ 4.3.2 Expected solar-like oscillations within PLATO P5 sample
Let us note that 10% of the P5 sample will be observed with a 50 s cadence. This will make it possible to detect solar-like oscillations and to measure at least the global seismic parameters and . We performed the same probability calculation as for the P1–P2 sample after eliminating the same types of stars and assuming a detection probability equal or greater than 99% (and ). The number of seismic positive detections is estimated to be 9941 stars for an observation period of 2 years. This number falls to 5637 stars after only 1 year of observation ad only 401 stars after 30 days of observations. In the last case, none are expected to be in the main sequence: the sample is dominated by subgiants because their amplitudes (roughly ) are higher than for main-sequence stars. The drastic increase of detections of oscillating stars with the observing time in the P5 sample is illustrated in Fig. 12.
@@ 4.4 PLATO seismic performance for stellar mass, radius, and age characterisation for the P1–P2 sample
For the subset of P1–P2 stars with expected solar-like oscillations, the detection and highly precise measurement of individual frequencies for a significant number of modes is ensured by the selection of a high signal-to-noise ratio by construction. This will allow the determination of the stellar age at the level of 10% for F and G stars. As an illustration of the expected PLATO performance, [268], to which we refer for details, illustrated the excellent PLATO performance for the Kepler legacy star, 16 Cyg B, a 6th magnitude dwarf observed over 815 days. The simulated PLATO power spectrum built assuming the expected noise for a star observed with 24 cameras at the end-of-life (EOL) conditions and for 2 years of observations with the PSLS simulator reveals the strong capacity of the mission to detect the oscillations (note that the noise level was for 28 telescopes (34 ppm h1/2) at the time).
@@
Dipole () modes produce the highest amplitude and smallest uncertainty, so we focus on these here. Denoting the uncertainty of the mode closest to as , Kepler data taught us that a precision on frequencies of the order of = 0.2Hz for a few modes around the frequency at maximum power can provide an age precision at the level of 10% for a Sun-like star. Several hare-and-hound exercises using artificial data constructed for the PLATO noise characteristics were conducted by PMC members (e.g. Goupil 130, Cunha et al 94) and confirmed this result. The artificial star named Zebedee and studied by [94] has a mass, radius, and effective temperature similar to the PLATO reference star and is situated close to the PIC star, PICid 8525275 ( mag). With an assumed PLATO noise level for a 9th magnitude star, the frequency uncertainty was found to be = 0.13Hz at and the age uncertainty was 13%. The target was estimated to be 3.085 Gyr old, i.e., younger than the Sun. Rescaling to the solar age, the age uncertainty is at the level of 8.8%. When the Zebedee data are degraded to the level that the frequency uncertainty increases to Hz, corresponding to a mag star, [94] found an age uncertainty of the order of 15%. This becomes 10 % when rescaling to the Sun’s age. Another practical yet more conservative estimation results from using [15], leading to a relative age uncertainty of 17% for Zebedee, turning into a relative age uncertainty of when rescaling to the Sun’s age at 10.5 mag.
[198] computed how the frequency uncertainties for individual modes decrease with increasing observation duration. A light curve was simulated for a star of mag with . The foreseen PLATO reference noise level of at magnitude 11 for 28 telescopes at the time of this study was used. With an updated current reference noise level of at magnitude 11 (Börner et al. in prep.), the frequency uncertainties correspond to a simulated star of mag. The frequency uncertainty increases with increasing magnitude but decreases with decreasing effective temperature. In this way, it is found that a level of Hz can be reached after more than roughly 1.5 years of observations. We will then take that criterion to estimate the number of stars for which one can reach a statistical error of 10% for the stellar age of a reference star.
Assuming individual frequencies are available, we determine the uncertainty on the frequencies using the [191] formula, which depends on the total SNR and the duration of the observation and has been proven to yield the right order of magnitude. Figure. 13 shows the histogram of the frequency uncertainty computed from the formulation by [191] for each P1–P2 star for which the detection probability is equal to or larger than 99% and with masses less or equal to 1.2. The bulk of stars have frequency uncertainties below about 0.12 Hz. Those estimates are purely theoretical but give the same order of magnitude than for the Kepler Legacy sample.
@@
We now turn to the expected accuracy of the seismic determinations of masses, radii, and ages. For the mass and radius relative uncertainties, we used empirical formulations obtained as fits of the results of seismic inferences of mass, radius and age of synthetic stellar models (see Goupil et al. in prep. for detail). For the age uncertainty, as mentioned above, we consider the criterion Hz as a proxy for the requirement on the age uncertainty (i.e. than 10%). Those estimates indicate that roughly 5271 MS-stars will have mass and radius uncertainties below 15% and 2% respectively. Among them about 2651 stars will have a frequency uncertainty Hz. Figures 14, and 15 focus on MS-stars with masses and show the , and radius histograms for stars satisfying the above constraints : mass and radius relative uncertainties better than 15 % and 2% respectively and Hz . Of interest for the exoplanet yields, 1640 stars with masses and 937 stars with radius have Hz.
@@ 4.5 Expectations for M dwarfs (P4 Sample)
In the framework of the PLATO project, an updated and accurate preliminary library of stellar models for M-dwarfs – stars with – has been produced. This model set is based on the best available input physics as far as opacity and thermodynamical tabulations, nuclear cross sections, and outer boundary conditions are concerned. Comparison with suitable observational benchmarks has shown good agreement between model predictions and observations, although some discrepancies occur. More precisely, in the regime of fully convective stars (), stellar models over-predict the effective temperature of suitable benchmark stars by on average 3—4%, and underestimate the stellar radius by about 5%. Current theoretical mass – IR-band magnitude luminosity relations to derive the mass of M dwarfs deliver an accuracy of the order of 4% (e.g., Passegger et al 245). Quantifying the uncertainty of the age estimation is, however, quite difficult due to the intrinsic extremely long evolutionary lifetimes of M dwarfs. A search for suitable binary systems formed by an M dwarf and a more massive star is ongoing within the PMC. Improvements in M-dwarf modelling are expected on the topics of better treatment of the outer boundary conditions – a critical issue in the computation of M dwarf models – of convection and its link with the presence of magnetic fields and with the process of formation via mass accretion during the pre-main-sequence stage. All these poorly known phenomena affect the structural and evolutionary properties of these intriguing low-mass host stars.
@@ 4.6 Measurement of surface rotation period, stellar activity indices and flares
All the PLATO photometric timeseries will be analysed to search for surface rotation periods, surface differential rotation, and magnetic indicators and activity cycles. The analyses will be carried out on the data collected during each quarter, and on consecutive stitched quarters as the mission progresses. Such techniques are very sensitive to contamination from instrumental modulations. The long-term stability of the instrument at low-frequency is therefore crucial as discussed by [270] for instance. The time series will be analysed using a combination of different algorithms, such as the Generalised Lomb-Scargle (GLS; Handberg and Lund 140 and references therein) and the AutoCorrelation Function (ACF; McQuillan et al 209, Garcia et al 122), to produce a composite spectrum (hereafter CS) [66, 67]. Only highly-significant periodicities in the GLS power spectrum will be considered. Once the rotation period is identified in the CS, a machine learning (ML) random-forest approach is used to decide if the final rotation period will be the one computed from the GLS, the ACF or the CS (see, e.g., the Random fOrest Over STEllar Rotation (ROOSTER) methodology; [51, 271]). Some additional stellar parameters such as luminosity or effective temperature can also be used to help in such selection.
@@
An important point of such a methodology is that a random forest allows us to use more parameters in the decision than a manually-set threshold when assessing the robustness of a rotation detection and selecting the corresponding rotation period. This methodology has the advantage of preserving the uncertainty and the posterior distribution of the method selected by the random forest. This type of machine learning algorithm can be combined with any inference techniques applied on the light curve or the periodogram. In the baseline version of the PLATO pipeline, the uncertainty associated with the rotation period and the activity cycle length will be inferred from the width of the corresponding power spectrum peak, and it is expected to improve as the timeseries get longer. Candidate values for starspots and levels of differential rotation will be inferred from the relative difference between the rotation period and other, if any, significant periodicities. Such candidate values will be validated using priors from theoretical models, as well as information on either the single or binary nature of the star, since unresolved secondary components can also produce detectable periodicities in the flux timeseries. The presence and length of long-term periodicities, likely arising from either activity cycles or from beating of close frequencies, will be inferred as done for the rotation period, but focusing on the lower-frequency region of the power spectrum computed from the stitched timeseries. The baseline version of the PLATO pipeline may miss the detection of rotational modulation signals in stars with a low level of activity and active regions with lifetimes shorter than the rotation period such as the Sun during some specific activity phases along its activity cycle (cf. Reinhold et al 256). In such cases, we plan to apply other methods to estimate the stellar rotation period in future versions of the PLATO pipeline. Although with a lower precision of about 15-20%, those methods can be applied to stars with a low-activity level at an age comparable or older than the Sun (e.g. Amazo-Gómez et al 11, Reinhold et al 257).
@@
5 The PLATO Complementary Science programs
5.1 Background
PLATO’s Complementary Science (PLATO-CS hereafter) is dedicated to exploit the mission’s capacity for answering scientific questions outside its core remit and can be addressed via successful proposals within PLATO’s Guest Observer (GO) programme (see Section 10). Up to 8% of PLATO’s science data rate will be spent on PLATO-CS and will make it a unique legacy programme.
Previous photometric space missions built for exoplanet detection by the transit method, such as CoRoT [18], Kepler [179], and TESS [259] have already shown that the uninterrupted high-precision photometric light curves they assembled triggered studies far beyond the nominal mission goals. PLATO-CS will combine the best of Kepler and TESS in that it will provide a legacy programme based on long-term monitoring capacity for a huge sky coverage per Field-of-View (FoV), containing stars with a dynamical range in brightness covering ten orders of magnitudes, for sampling rates of 2.5 s, 25 s, 50 s, and 10 min. PLATO-CS will be able to rely on all raw (L0) and calibrated (L1) PLATO light curves, as well as imagettes for tens of thousands of them.
5.2 Support for the community
The PLATO-CS activity within ESA’s Science Working Team and supported by the PMC has several operational aims. First of all, it will help the worldwide community to prepare competitive GO programs (see Section 10) by means of simulated data and by using Kepler and TESS as preparatory tools and testbeds. The PLATO simulation PlatoSim [162] offers a versatile tool for the community to help prepare the assessment of PLATO’s capacity for various scientific goals. PLATO-CS will encourage and trigger GO proposals covering a variety of topics in Solar System science, stellar astrophysics, the Milky Way galaxy, transients and more broadly extragalactic science. Although we anticipate a large over-subscription factor for the GO call, the PLATO-CS team will facilitate relevant simulations with PlatoSim to showcase PLATO’s capacity for the CS, which will stimulate interest in the worldwide community.

### v2

@@ 4.3.1 Expected solar-like oscillations within the PLATO P1-P2 sample
The first question is to what extent solar-like oscillations can be detected for PLATO targets. Figure 11 shows the stars in the P1–P2 sample for two long-pointing fields in an HR diagram. They are taken from the PIC version PICv1.1.0 [252], which provides effective temperatures and stellar radii, hence luminosities. In Fig. 11, we distinguish between stars for which a theoretical calculation has led to a probability of detection of solar-type oscillations and those for which the threshold for a positive seismic detection has not been reached. The probability of detection were not calculated for hot stars in the instability band defined according to the criterion adopted by [75]. This concerns only a few stars because the restriction on the temperature of the hot side adopted to construct the PIC is more severe. We also did not consider the early red giants –also contained in the PIC– which will be included in a specific scientific calibration sample catalogue as part of the overall PIC [5]. Their performance will be addressed elsewhere. Once early red giants and hot stars have been removed, the total subsample (dwarfs and subgiants) contains 14083 stars. We considered a positive seismic detection when the probability of the signal being due to noise is below 0.1% and the probability of detecting solar-like oscillations is equal to 99% or above. Details of the computation can be found in [142].
The seismic detection probability depends on the observing duration. We therefore show two cases in Fig.11 : observing runs of 30 days and 2 years long. The probability also involves the signal-to-noise ratio in a power spectrum. We then used a combination of the formulations by [75] and by [291] for the oscillation amplitudes calibrated with Kepler data (a compilation of oscillating stars from the catalogues of [302] and [227] and short-cadence stars found with no detection from the [226] catalogue. For the noise, we used the PLATO (random and systematic residuals) noise level included in the PIC, to which we added the stellar granulation background noise. For comparison, we included in the plot the positions of some Kepler stars from the Legacy sample [218, 306] as characterised by [95].
The calculation of the probability to detect oscillations depends on the width () of the assumed Gaussian-shape envelope due to oscillations in a power spectrum. From Kepler observations, ranges from to . The first option ( together with a probability threshold set at 0.90) (later option 1) provides a number of positive detections in agreement with the number of Kepler stars with detected oscillations, but a too large number of false-positive detections among the Kepler stars with no detected oscillations. For the other option ( together with a conservative probability threshold of 0.99) (later on option 2), it is the reverse, i.e., the number of predicted non-detections is in agreement with the number of Kepler stars without oscillations, but too few detections are found compared to the Kepler stars with detected oscillations. Unless indicated otherwise, we remain conservative and use the second option to derive the number of expected detections for PLATO.
Tab. 3 shows the significant impact on the number of oscillation detection of the observing time (30 days and 2 years) assuming option 2 for different subsamples of the P1P2 sample. The figures are limited to the case of MS-stars with masses .
Had we assumed option 1 for the width of the oscillation envelope, the number of expected detections for stars with mass for instance would increase from about 55% of stars (option 2) up 76% of stars (option 1) for a 2 years run.
| cases | 730 days | 30 days |
| Detections | ||
| MS () + | ||
| subgiants | 11109 | 3188 |
| MS () | 4972 | 477 |
| subgiants | 6137 | 2711 |
| MS + subgiants | ||
| () | 2344 | 302 |
| MS () | 1844 | 166 |
| Perfomances ( MS with ) | ||
| 1844 | 166 | |
| 1844 | 153 | |
| and Hz | ||
| 1600 | 0 | |
| and Hz |
Not surprisingly, stars for which one might not detect solar-like oscillations for too short an observing time are MS-stars of low mass because their oscillation amplitudes are too small. In contrast, detection of solar-like oscillations - when applied to the Kepler sample- is not obtained for some stars for which they are theoretically expected [75, 74, 226]. Several reasons may occur, one being the strong magnetic activity. As an example, we found 8 % false-positive detections for the [226] Kepler sample. Based on this, the number of PLATO detections for main-sequence stars with mass decreases to 51%.
@@ 4.3.2 Expected solar-like oscillations within PLATO P5 sample
Let us note that 10% of the P5 sample will be observed with a 50 s cadence. This will make it possible to detect solar-like oscillations and to measure at least the global seismic parameters and . We performed the same probability calculation as for the P1–P2 sample after eliminating the same types of stars and assuming a detection probability equal or greater than 99% (and ). The number of seismic positive detections is estimated to be 9941 stars for an observation period of 2 years. This number falls to 5637 stars after only 1 year of observation and only 401 stars after 30 days of observations. In the last case, none are expected to be in the main sequence: the sample is dominated by subgiants because their amplitudes (roughly ) are higher than for main-sequence stars. The drastic increase of detections of oscillating stars with the observing time in the P5 sample is illustrated in Fig. 12.
@@ 4.4 PLATO seismic performance for stellar mass, radius, and age characterisation for the P1–P2 sample
For the subset of P1–P2 stars with expected solar-like oscillations, the detection and highly precise measurement of individual frequencies for a significant number of modes is ensured by the selection of a high signal-to-noise ratio by construction. As an illustration of the expected PLATO performance, [291], to which we refer for details, illustrated the excellent PLATO performance for the Kepler legacy star, 16 Cyg B, a 6th magnitude dwarf observed over 815 days. The simulated PLATO power spectrum built assuming the expected noise for a star observed with 22 cameras at the end-of-life (EOL, see Section 11) conditions and for 2 years of observations with the PSLS simulator reveals the strong capacity of the mission to detect the oscillations (note that the noise level was for 28 telescopes (34 ppm h1/2) at the time).
This will allow the determination of the stellar age at the level of 10% for F and G stars. To give a general estimate of the PLATO stellar performances in terms of stellar ages is difficult, here we use a proxy based on the expected frequency uncertainties.
@@
Dipole () modes produce the highest amplitude and smallest uncertainty, so we focus on these here. Denoting the uncertainty of the mode closest to as , Kepler data taught us that a precision on frequencies of the order of = 0.2Hz for a few modes around the frequency at maximum power can provide an age precision at the level of 10% for a Sun-like star. Several hare-and-hound exercises using artificial data constructed for the PLATO noise characteristics were conducted by PMC members (e.g. Goupil 141, Cunha et al 99) and confirmed this result.
Another practical yet more conservative estimation results from using [16].
[218] computed how the frequency uncertainties for individual modes decrease with increasing observation duration. A light curve was simulated for a star of mag with . The foreseen PLATO reference noise level of at magnitude 11 for 28 telescopes at the time of this study was used. With an updated current reference noise level of at magnitude 11 [47], the frequency uncertainties correspond to a simulated star of mag. The frequency uncertainty increases with increasing magnitude but decreases with decreasing effective temperature. In this way, it is found that a level of Hz can be reached after more than roughly 1.5 years of observations. We will then take that criterion to estimate the number of stars for which one can reach a statistical error of 10% for the stellar age of a reference star.
Assuming individual frequencies are available, we determine the uncertainty on the frequencies using the [209] formula, which depends on the total SNR and the duration of the observation and has been proven to yield the right order of magnitude [we refer to the monograph by 27, for thorough discussions on the basic proportionality with for the frequency uncertainty occurring in this formula].
Figure. 13 shows the histogram of the frequency uncertainty computed from the formulation by [209] for each P1–P2 star for which the detection probability is equal to or larger than 99% and with masses less or equal to 1.2. The bulk of stars have frequency uncertainties below about 0.12 Hz. Those estimates are purely theoretical but give the same order of magnitude than for the Kepler Legacy sample.
@@
We now turn to the expected accuracy of the seismic determinations of masses, radii, and ages. For the mass and radius relative uncertainties, we used empirical formulations obtained as fits of the results of seismic inferences of mass, radius and age of synthetic stellar models (see [142] for detail). For the age uncertainty, as mentioned above, we consider the criterion Hz as a proxy for the requirement on the age uncertainty (i.e. than 10%). Those estimates are given in Tab. 3 above. Figures 14 and 15 focus on MS-stars with masses and show the , and radius histograms for stars satisfying the above constraints : mass and radius relative uncertainties better than 15 % and 2% respectively and Hz. Of interest for the exoplanet yields, 1640 stars with masses and 937 stars with radius have Hz.
@@ 4.5 Expectations for M dwarfs (P4 Sample)
In the framework of the PLATO project, an updated and accurate preliminary library of stellar models for M-dwarfs - stars with - has been produced. A sub-set of this library has been already published (e.g. [161, 274], but for the aims of the PLATO project, the set has been hugely extended by adopting a very fine grid as far as it concerns the total stellar mass and metallicity (e.g. Orlander et al. 2024, in preparation). All those models are available to the scientific community.
This model set is based on the best available input physics as far as opacity and thermodynamical tabulations, nuclear cross sections, and outer boundary conditions are concerned.
Comparison with suitable observational benchmarks has shown the existence of a good agreement between model predictions and observations, although some discrepancies occur. More in detail, in the regime of fully convective stars (), stellar models over-predict the effective temperature of suitable benchmark stars by on average 3—4%, and underestimate the stellar radius by about 5%. Current theoretical mass – IR-band magnitude luminosity relations to derive the mass of M dwarfs deliver an accuracy of the order of 4% (e.g., Passegger et al 268). The presence of these discrepancies is a clear proof that there is still a significant uncertainty in some input physics adopted in the stellar modelling, and/or some physical process such as magnetic field/rotation is not properly accounted in the computation of models for such stars. In order to solve this problem and crucially improve our capability to characterise host-planet M dwarfs, a detailed search for well studied single M dwarfs and, in particular, suitable binary systems formed by an M dwarf and a more massive star is ongoing within the PMC. The selection of these crucial benchmark M dwarf stars will allow us in the context of the PLATO stellar science, to achieve a better knowledge of the atmospheric layers of these stars, and hence of the outer boundary conditions needed in their modelling, as well as of the link between convection and magnetic fields. In this specific context, the availability of a detailed characterization of the magnetic field activity in these stars - obtained by combining the PLATO data analysis with the results collected from related follow-up surveys - is mandatory to shed light on the origin and physical properties of the magnetic fields in M dwarfs and their impact on their structural properties. On a different ground is the issue of a realistic estimation of the uncertainty on the age estimate for these stars; this is made very difficult by the intrinsic extremely long evolutionary lifetime of these peculiar stars.
@@ 4.6 Measurement of surface rotation period, stellar activity indices and flares
All the PLATO photometric timeseries will be analysed to search for surface rotation periods, surface differential rotation, and magnetic indicators and activity cycles. The analyses will be carried out on the data collected during each quarter, and on consecutive stitched quarters as the mission progresses. Such techniques are very sensitive to contamination from instrumental modulations. The long-term stability of the instrument at low-frequency is therefore crucial as discussed by [293] for instance. The time series will be analysed using a combination of different algorithms, such as the Generalised Lomb-Scargle (GLS; Handberg and Lund 153 and references therein) and the AutoCorrelation Function (ACF; McQuillan et al 231, Garcia et al 133), to produce a composite spectrum (hereafter CS) [72, 73]. Only highly-significant periodicities in the GLS power spectrum will be considered. Once the rotation period is identified in the CS, a machine learning (ML) random-forest approach is used to decide if the final rotation period will be the one computed from the GLS, the ACF or the CS (see, e.g., the Random fOrest Over STEllar Rotation (ROOSTER) methodology; [55, 294]). The light curves included in the training set will be carefully selected, with the possibility to incorporate relevant light curves from previous space missions (Kepler/K2, TESS) together with PLATO simulations. It is planned that the module performance will be evaluated within the PLATO consortium after launch, once the first observations are available, and the possibility to include a subset of real PLATO targets with rotation measurements validated independently from ROOSTER is under consideration [56]. Some additional stellar parameters such as luminosity or effective temperature can also be used to help in the selection of the best rotation period.
@@
An important point of such a methodology is that a random forest allows us to use more parameters in the decision than a manually-set threshold when assessing the robustness of a rotation detection and selecting the corresponding rotation period. This methodology has the advantage of preserving the uncertainty and the posterior distribution of the method selected by the random forest. This type of machine learning algorithm can be combined with any inference techniques applied on the light curve or the periodogram. In the baseline version of the PLATO pipeline, the uncertainty associated with the rotation period and the activity cycle length will be inferred from the width of the corresponding power spectrum peak, and it is expected to improve as the timeseries get longer. Candidate values for starspots and levels of differential rotation will be inferred from the relative difference between the rotation period and other, if any, significant periodicities. Such candidate values will be validated using priors from theoretical models, as well as information on either the single or binary nature of the star, since unresolved secondary components can also produce detectable periodicities in the flux timeseries. On the other hand, starspot evolution and differential rotation may affect the determination of the mean stellar rotation period for some targets. In most solar-like stars, we expect an increase in the broadening of the periodogram peaks well within our adopted 10% uncertainty. However, a systematic investigations of these effects for a proper consideration in PLATO data analysis is under way (see [56]). The presence and length of long-term periodicities, likely arising from either activity cycles or from beating of close frequencies, will be inferred as done for the rotation period, but focusing on the lower-frequency region of the power spectrum computed from the stitched timeseries.
In the future, some complementary methods will be evaluated in order to be included along the current baseline. Among possible additions, we mention the wavelet decomposition (Torrence and Compo 319, Liu et al 216, Mathur et al 225), a time-frequency analysis that eases discriminating stellar rotation signatures with instrumental modulations or neighbours contamination. A new development around this wavelet approach, the Gradient Power Spectrum (GPS) was recently proposed in order to measure rotation in stars with low levels of activity (e.g. Amazo-Gómez et al 12, Reinhold et al 280).
@@
5 The PLATO Complementary Science
Previous photometric space missions built for exoplanet detection by the transit method, such as CoRoT [19], Kepler [192], and TESS [282] have already shown that the uninterrupted high-precision photometric light curves they assembled triggered studies far beyond the nominal mission goals. Following this, the PLATO Complementary Science programme (PLATO-CS hereafter) stands for a broad activity initiated by ESA’s Science Working Team and supported by the PMC. It has the global goal to get the maximum scientific return from the PLATO mission. In order to achieve this, PLATO-CS has defined several operational tasks.